A balloon is filled with 12 L of air at a pressure of 2 atm. What is the volume of the balloon if the pressure is changed to 3 atm?
1
Expert's answer
2020-04-23T13:07:54-0400

p1V1 = p2V2

V2 = p2V2/p2 = 2*12/3 = 8 L